OpenStudy (anonymous):

1/p + 1/q = 1/f. How do you solve for f?

OpenStudy (anonymous):

\[ f = \frac{1}{\frac{1}{p}+\frac{1}{q}} = \frac{pq}{p+q} \]

OpenStudy (anonymous):

How do you get there?

OpenStudy (anonymous):

Just take the reciprocal on both sides.

OpenStudy (anonymous):

I still don't understsand the steps.. I'm very new to this.

OpenStudy (jamesj):

1/p + 1/q = (p + q)/pq. Hence if 1/f = 1/p + 1/q, that means 1/f = (p + q)/pq. Taking the reciprocal of both sides, f = pq/(p + q).